OpenStudy (anonymous):

find \(n\) if \(n=\sum _{k=1}^nk\)

OpenStudy (anonymous):

n=1 works, doesn't it? \[\sum_{k=1}^{1}k = 1\]

OpenStudy (anonymous):

yes it does ,is that the only case

OpenStudy (kirbykirby):

I would think so. Can you imagine a sum 1+2+...+n = n, if n is other than 1?

OpenStudy (anonymous):

yeah...1+2 > 2 1 + 2 + 3 > 3 .....the sum is always gonna be bigger than n for the rest of the integers

ganeshie8 (ganeshie8):

n = n(n+1)/2

OpenStudy (anonymous):

the question further asks us to generalise \(pk=\sum_{k=1}^nk\)

OpenStudy (anonymous):

\(p\) is a positive integer

OpenStudy (anonymous):

i meant \[nk=\sum_{k=1}^nk\]according to @ganeshie8 set up \(pn=\dfrac{n(n+1)}{2}\) \(2pn=n^2+n\implies n^2+n(1-2p)\) \[n=2p-1\] so this holds for odd numbers

ganeshie8 (ganeshie8):

u meant this right : \(np=\sum \limits_{k=1}^nk\)

OpenStudy (anonymous):

yes yes thank you @ganeshie8 @kirbykirby @BangkokGarrett

ganeshie8 (ganeshie8):

oh nice :) looking at \(n = n(n+1)/2\), n = 0 doesnt work since the lower bound is 1 is it ?

ganeshie8 (ganeshie8):

for ur first part of q..

OpenStudy (anonymous):

hmmmm yes ,actually it wpuld work for \(np\) case as well

OpenStudy (anonymous):

i mean it doesnt work for both cases

ganeshie8 (ganeshie8):

Got it !! ty :)